# Puzzle: Distillation-Based NAS for Inference-Optimized LLMs

### v4

@@ Abstract
We showcase our framework’s impact via Llama-3.1-Nemotron-51B-Instruct (Nemotron-51B), a publicly available model derived from Llama-3.1-70B-Instruct. Nemotron-51B achieves a 2.17$\times$ inference throughput speedup, fitting on a single NVIDIA H100 GPU while retaining 98.4% of the original model’s benchmark accuracies. Notably, it is the most accurate model supporting single H100 GPU inference with large batch sizes, despite training on only 45B tokens, far fewer than the 15T used to train Llama-70B. Lastly, we derive Llama-3.3-Nemotron-49B-Super-Base to demonstrate Puzzle can retain long-context and that lightweight alignment on these derived models allows them to surpass the parent model in specific capabilities. Our work establishes that powerful LLM models can be optimized for efficient deployment with only negligible loss in quality, underscoring that inference performance, not parameter count alone, should guide model selection.
@@ 1 Introduction
Our framework offers several advantages. First, it enjoys extremely low costs relative to training a model from scratch. For instance, the entire training—BLD before the MIP stage and GKD afterward—required less than 50B tokens to run on Llama-3.1-70B-Instruct, compared to more than 15T tokens used to train the parent model. Additionally, our method requires only the parent model’s weights—not its training data—making it ideal for “open-weights, closed-data” scenarios where training data of the parent model is not publicly available. This allows practitioners to take any freely-available model and tailor it to their specific hardware or use case. To demonstrate the effectiveness of our framework, we present Llama-3.1-Nemotron-51B-Instruct (Nemotron-51B), derived from the Llama-3.1-70B-Instruct parent model using Puzzle. Nemotron-51B breaks the efficient frontier of LLMs on a single NVIDIA H100 GPU, establishing a new state-of-the-art for commercial applications, by achieving unmatched throughput and memory efficiency on this hardware. Interestingly, the Nemotron-51B resulting architecture is unique and irregular, with many layers featuring reduced or skipped attention and FFN operations. This design enhances NVIDIA H100 utilization under FP8 quantization while preserving accuracy. We also introduce a derivative of Llama-3.1-8B-Instruct, which breaks the efficient frontier for its throughput slice. While Nemotron-51B also leads within its parameter range, we argue that categorizing models solely by parameter count—such as 50B or 70B—is inadequate for real-world applications. Instead, inference performance under specific hardware, inference engine, quantization levels, budget constraints, and usage profiles—such as varying sequence lengths and batch sizes—should guide model selection.
@@
(2) Using Puzzle, we introduce Llama-3.1-Nemotron-51B-Instruct, a model optimized for the NVIDIA H100 GPU thus setting a new benchmark for throughput and memory efficiency in commercial applications. Our results demonstrate that hardware-specific architecture optimization can achieve superior efficiency compared to traditional uniform architectures while maintaining model performance.
@@ 4.2 Scoring Architecture Solutions
We consider several metrics as potential replace-1-block scores: (1) downstream accuracy: accuracy on downstream tasks such as MMLU [20]. While downstream accuracy is popular in Computer Vision applications, it can be very costly to measure in LLMs, especially given the number of block variants requiring scoring. (2) LM loss: causal language modeling loss, defined as the average log-likelihood of a validation corpus under the model’s next-token prediction distribution. (3) KL divergence: Kullback–Leibler divergence between the next-token prediction distributions of the evaluated model and the parent model, averaged across tokens in a validation corpus. KL divergence is widely used in KD setups as a statistical distance measure but has not been explored for decomposed NAS scoring until now. Our analysis in Section 8.1.4 highlights its effectiveness for block scoring.
@@ 7 Main Results
Throughput comparison: Table 3 specifies the throughput performance of Nemotron-51B against its parent model across diverse input-output sequence lengths. Nemotron-51B achieves speedups up to 2.17x, enabling larger workloads per GPU and making it highly efficient for deployment.
Table 3: Throughput comparison of Nemotron-51B and Llama-3.1-70B-Instruct across various scenarios. Throughput is measured in tokens per second per GPU (NVIDIA H100). TP# indicates the number of GPUs used in tensor parallelism. Note: Results were obtained on NVIDIA H100 SXM GPUs with FP8 quantization for weights, activations and KV cache using TensorRT-LLM. Optimal tensor parallelism was used for each model. Input/output sequence lengths indicate the prefill (input) and decode (output) operations performed by the LLM.
@@
To demonstrate this extension, we produced Llama-3.3-Nemotron-49B-Super-Base (henceforth Nemotron-49B-Base, the base version of Llama-3.3-Nemotron-49B-Superhttps://huggingface.co/nvidia/Llama-3_3-Nemotron-Super-49B-v1) from Llama-3.3-70B-Instruct with identical constraints to Nemotron-51B and uptrained on an additional 5B tokens at 64K context and 5B at 128K. Table 4 shows that Nemotron-49B-Base maintains or exceeds its parent’s performance up to 16K tokens, retains over 98% at 64K, and remains above 94% at 128K. Full performance details are provided in Appendix B.
@@ 9 Conclusions and Future Directions
In this work we present Puzzle, a framework that modifies LLMs from their over-parameterized training configurations to optimized, inference-efficient architectures tailored for specific hardware. Our results challenge the conventional wisdom that model architectures should remain uniform across training and inference, demonstrating that post-training architectural optimization can significantly improve efficiency without compromising performance. Using Puzzle, we developed Nemotron-51B from the Llama-3.1-70B-Instruct model, achieving unmatched throughput and memory efficiency on a single NVIDIA H100 GPU. Nemotron-51B maintains over 98% of the parent model’s accuracy while delivering a 2.2x throughput improvement in optimized settings and up to 5x improvement on a single GPU, demonstrating the effectiveness of optimizing for deployment environments rather than focusing on model size alone.
The Puzzle framework achieves these improvements with remarkable efficiency in training resources. Requiring fewer than 50B tokens—compared to the trillions needed to train models from scratch—Puzzle produces high-performing models at a fraction of the usual cost. This extreme search and training efficiency still results in drastic inference performance improvements of Puzzle optimized models, thus enabling widespread and efficient deployment of LLMs across data centers and edge devices.
The success of Puzzle opens several promising directions for future research. Our introduction of decoupled BLD, which is significantly more efficient than coupled BLD, makes it feasible to evaluate a much larger set of potential blocks within a single optimization run. This efficiency enables exploration of novel operations as alternative Puzzle blocks, such as variable window attention mechanisms [9], state-space models [16, 15], or other architectural innovations. The framework could also be extended to optimize models for specific capabilities, such as Chain-of-Thought reasoning or multimodal tasks, including vision-language models [34, 57] and retrieval-augmented generation [30].
Future work might also explore more sophisticated search algorithms beyond MIP, incorporating reinforcement learning or evolutionary approaches to discover even more efficient architectures. The relationship between architectural patterns and specific model capabilities deserves deeper investigation, as understanding this connection could lead to more targeted optimization strategies. Additionally, studying the robustness of Puzzle-optimized models to distribution shifts and their ability to adapt to new tasks through fine-tuning would be valuable, particularly in understanding how architectural modifications affect model reliability and exploring methods for dynamic architecture adaptation.
Puzzle represents a significant step toward making advanced language models more practical and accessible. The framework’s ability to customize state-of-the-art models based on hardware constraints, while maintaining high performance, points the way toward a future where powerful AI systems can be affordably deployed across a wide range of environments and use cases, advancing the field of generative AI in a cost-effective and application-oriented direction.

### v5

@@ Abstract
We showcase our framework’s impact via Llama-3.1-Nemotron-51B-Instruct (Nemotron-51B) and Llama-3.3-Nemotron-49B, two publicly available models derived from Llama-70B-Instruct. Both models achieve a 2.17$\times$ inference throughput speedup, fitting on a single NVIDIA H100 GPU while retaining 98.4% of the original model’s benchmark accuracies. These are the most accurate models supporting single H100 GPU inference with large batch sizes, despite training on 45B tokens at most, far fewer than the 15T used to train Llama-70B. Lastly, we show that lightweight alignment on these derived models allows them to surpass the parent model in specific capabilities. Our work establishes that powerful LLM models can be optimized for efficient deployment with only negligible loss in quality, underscoring that inference performance, not parameter count alone, should guide model selection.
@@ 1 Introduction
Our framework offers several advantages. First, it enjoys extremely low costs relative to training a model from scratch. For instance, the entire training—BLD before the MIP stage and GKD afterward—required less than 50B tokens to run on Llama-3.1-70B-Instruct, compared to more than 15T tokens used to train the parent model. Even smaller budgets are possible—down to 4B tokens—while still preserving strong performance (see Table 7). Additionally, our method requires only the parent model’s weights—not its training data—making it ideal for “open-weights, closed-data” scenarios where training data of the parent model is not publicly available. This allows practitioners to take any freely-available model and tailor it to their specific hardware or use case. To demonstrate the effectiveness of our framework, we present Llama-3.1-Nemotron-51B-Instruct (Nemotron-51B), derived from the Llama-3.1-70B-Instruct parent model using Puzzle. Nemotron-51B breaks the efficient frontier of LLMs on a single NVIDIA H100 GPU, establishing a new state-of-the-art for commercial applications, by achieving unmatched throughput and memory efficiency on this hardware. Interestingly, the Nemotron-51B resulting architecture is unique and irregular, with many layers featuring reduced or skipped attention and FFN operations. This design enhances NVIDIA H100 utilization under FP8 quantization while preserving accuracy. We also introduce a derivative of Llama-3.1-8B-Instruct, which breaks the efficient frontier for its throughput slice. While Nemotron-51B also leads within its parameter range, we argue that categorizing models solely by parameter count—such as 50B or 70B—is inadequate for real-world applications. Instead, inference performance under specific hardware, inference engine, quantization levels, budget constraints, and usage profiles—such as varying sequence lengths and batch sizes—should guide model selection.
@@
(2) Using Puzzle, we introduce Nemotron-51B and Nemotron-49B, both optimized for a single H100 GPU, thus setting a new benchmark for commercial applications. We show Nemotron-49B preserves its parent’s 128K context via short uptraining and benefits from RLHF alignment—outperforming Nemotron-51B on some tasks. Puzzle’s robustness is demonstrated throughout the paper by applying it numerous times with varied constraints, datasets, budgets and parent models.
@@ 4.2 Scoring Architecture Solutions
We consider several metrics as potential replace-1-block scores: (1) downstream accuracy: accuracy on downstream tasks such as MMLU [22]. While downstream accuracy is popular in Computer Vision applications, it can be very costly to measure in LLMs, especially given the number of block variants requiring scoring. (2) LM loss: causal language modeling loss, defined as the average log-likelihood of a validation corpus under the model’s next-token prediction distribution. (3) KL divergence: Kullback–Leibler divergence between the next-token prediction distributions of the evaluated model and the parent model, averaged across tokens in a validation corpus. KL divergence is widely used in KD setups as a statistical distance measure but has not been explored for decomposed NAS scoring until now. Our analysis in Section 8.1.4 highlights the KL divergence effectiveness for block scoring.
@@ 7 Main Results
Throughput comparison: Table 3 specifies the throughput performance of Nemotron-51B against its parent model across diverse input-output sequence lengths. Nemotron-51B achieves speedups up to 2.17x, enabling larger workloads per GPU and making it highly efficient for deployment. For each model and hardware configuration, we automatically selected the optimal TP and batch size to maximize throughput per GPU. The inference engine handled this selection dynamically for each run. For example, Nemotron-51B achieved optimal throughput with TP=1 and batch size 256, while Llama-3.1-70B performed best with TP=4 and batch size 384.
Table 3: Throughput comparison of Nemotron-51B and Llama-3.1-70B-Instruct across various scenarios. Throughput is measured in tokens per second per GPU (NVIDIA H100). TP# indicates the number of GPUs used in tensor parallelism. Note: Results were obtained on NVIDIA H100 SXM GPUs with FP8 quantization for weights, activations and KV cache using TensorRT-LLM. Optimal tensor parallelism was used for each model. Input/output sequence lengths indicate the prefill (input) and decode (output) operations performed by the LLM. *TP=1 is not the optimal configuration for Llama-3.1-70B and is included for equal-resource comparison.
@@ 7.1 Additional Evaluations and Puzzle derivatives
To demonstrate this extension, we produced Llama-3.3-Nemotron-49B-Super-Base (henceforth Nemotron-49B-Base, the base version of Llama-3.3-Nemotron-Super-49Bhttps://huggingface.co/nvidia/Llama-3_3-Nemotron-Super-49B-v1) from Llama-3.3-70B-Instruct with identical constraints to Nemotron-51B and uptrained on an additional 5B tokens at 64K context and 5B at 128K. Table 4 shows that Nemotron-49B-Base maintains or exceeds its parent’s performance up to 16K tokens, retains over 98% at 64K, and remains above 94% at 128K. Full performance details are provided in Appendix B.
@@
Additional Puzzle derivatives: Beyond the aforementioned models, Puzzle was also used in the development of the following derivatives: (1) Puzzle was applied to Llama-3.1-405B-Instruct with constraints requiring a 1.5$\times$ latency speedup and compatibility with a single NVIDIA 8$\times$×H100 node (640 GB) or a single B100 GPU (192 GB). The resulting model retained 99.5% of the parent model’s accuracy (averaged across MMLU, MT-Bench, MMLU-Pro, HumanEval, and Arena Hard). FFN Fusion [10] was subsequently applied to further improve latency, followed by continued pretraining—resulting in Llama-3.1-Nemotron-Ultra-253B-CPT, which serves as the base for the publicly released Llama-3.1-Nemotron-Ultra-253B https://huggingface.co/nvidia/Llama-3_1-Nemotron-Ultra-253B-v1. (2) In Blakeman et al. [11], a variation of Puzzle—referred to as “miniPuzzle”—was applied to Nemotron-H-56B-Base under constraints targeting RTX 5090 deployment with a 1M context window. The resulting 47B model retained 99.94% of the parent model’s accuracy (averaged over MMLU, MMLU-Pro, GSM8K, and HellaSwag). The key distinction between miniPuzzle and Puzzle is that miniPuzzle enforces a homogeneous block design across all layers.
Training Token Budget: The GKD phase described earlier utilized 45B tokens, which might exceed practical budgets for some users. However, in practice, our method can achieve substantial accuracy recovery with significantly fewer tokens. Our experiments, summarized in Table 7, demonstrate that notable accuracy recovery can be realized with reduced token usage. After only 3.7B tokens of GKD, Nemotron-51B recovered 98.8% of its parent’s accuracy on MMLU and MT-Bench benchmarks. Similarly, Nemotron-49B regained 99.63% of its parent’s accuracy after only 8.68B tokens, and even 98.47% after just 2.9B tokens. These results indicate that while 45B tokens are modest compared to full model training from scratch, users can flexibly adjust GKD token counts based on available resources without severely compromising accuracy.
Table 7: Effectiveness of GKD training with reduced token budgets. The table shows the performance recovery relative to the parent model for Nemotron-51B and Nemotron-49B using various amounts of tokens in GKD.
Model GKD Tokens (B) Performance (MMLU / MT-Bench) Accuracy Preserved (%)
Nemotron-49B-Base 8.68 80.73 / 8.87 99.63
Nemotron-49B-Base 2.9 80.72 / 8.675 98.47
Nemotron-49B-Base 0.72 80.4 / 8.59 97.79
@@ 8.2 Search Algorithm Ablation Studies
  • Random Architecture Baselines: Nemotron-51B outperforms both fully random and random-from-library architectures under the same training and speed constraints, underscoring the role of the MIP solver in selecting effective architectural compositions (see Section 8.2.4).
@@ 8.2.3 Importance of Data-Driven Quality Estimation in Architecture Scoring
8.2.4 Evaluating Random Architectures as Search Baselines
To further validate the effectiveness of our search algorithm, we evaluated several baselines where block selection was performed randomly. All models were trained with 10B tokens.
The first baseline (“Random-from-block-library”) randomly samples block variants from the full block library while satisfying the same throughput constraint as Nemotron-51B, ignoring their block scores. The second (“Fully Random”) uses a completely random architecture that is not constrained to trained block variants, but adheres to the same speed constraints. Finally, we added a third baseline—Parent-Randomized—which evaluates Llama-3.1-70B with randomized weights but no architectural modifications.
As shown in Table 15, Puzzle significantly outperforms all baselines. Notably, the random-from-library baseline achieves only 86.6% of Nemotron-51B’s performance, despite being constructed from the same trained blocks. This experiment further emphasizes the role of MIP in selecting high-quality architectural compositions from the library.
Table 15: Comparison of Nemotron-51B with randomly constructed architectures, all trained for 10B tokens. The random-from-library variant uses trained blocks, while the fully random variant ignores the block library. The Parent-Randomized model uses Llama-70B’s architecture with random weights.
Model MMLU MT-Bench Avg. Accuracy Relative to Llama-70B
Nemotron-51B (10B tokens) 79.7 8.89 84.30 98.61%
Random-from-block-library 66.02 8.20 74.01 86.58%
Fully Random 23.13 0.89 16.02 18.73%
Parent-Randomized 23.42 0.95 16.46 19.25%
Llama-3.1-70B 81.66 8.93 85.48 100%
@@ 8.3 Global Knowledge Distillation Uptraining Importance
8.4 Comparison with Related Work
There are many different techniques to prune or compress models, most of which could be complementary to Puzzle. Some pruning methods, for example, could be used to construct alternative blocks for Puzzle’s block library. Nevertheless, in this section we compare several established methods directly against the basic Puzzle framework, without integrating them.
We compare Puzzle against structured sparsity (Wanda [53]) and low-rank approximation (similar to [30] with subsequent distillation) applied to Llama-3.1-70B under similar throughput constraints. Wanda applied 2:4 structured sparsity without additional training, while the low-rank method used factorized layers followed by distillation. Nemotron-51B significantly outperformed both methods, achieving 99.49% of the parent’s average accuracy (MMLU and MT-Bench), compared to 92.23% for Wanda and 88.96% for low-rank approximation (see Table 17). Subsequent distillation post-pruning with Wanda slightly improved MMLU (73.69) without impacting MT-Bench. Moreover, since both structured sparsity and low-rank approximations represent subsets of Puzzle’s broader search space, integrating these approaches into Puzzle could further enhance performance.
Table 17: Comparison of Puzzle, Wanda (structured sparsity), and low-rank approximation methods on Llama-3.1-70B derivatives under similar throughput constraints.
Model MMLU MT-Bench Average Accuracy Accuracy Preserved (%)
Nemotron-51B 80.20 8.99 85.05 99.49%
Wanda [53] 72.99 8.39 78.44 92.23%
Low-rank 72.87 8.01 76.05 88.96%
Llama-3.1-70B (Parent) 81.66 8.93 85.48 100%
Other methods such as Minitron [44], ShortGPT [41], and SlimGPT [37] share conceptual similarities with Puzzle, but each represents a constrained subset of Puzzle’s broader optimization space. For the most part, Minitron restricts modifications to homogeneous block replacements across all layers, ShortGPT focuses exclusively on redundant layer removal, and SlimGPT employs incremental pruning ratios via a fixed heuristic. Puzzle generalizes and extends these approaches, allowing for heterogeneous, layer-specific modifications, diverse block alternatives including no-op layers, and customizable pruning ratios optimized globally through MIP-based optimization.
@@ 9 Conclusions and Future Directions
In this work we present Puzzle, a framework that modifies LLMs from their over-parameterized training configurations to optimized, inference-efficient architectures tailored for specific hardware. Puzzle achieves these improvements with remarkable efficiency in training resources. Requiring fewer than 50B tokens—compared to the trillions needed to train models from scratch—Puzzle produces high-performing models at a fraction of the usual cost. This extreme search and training efficiency still results in drastic inference performance improvements of Puzzle optimized models, thus enabling widespread and efficient deployment of LLMs across data centers and edge devices.
The success of Puzzle opens several promising directions for future research. Our introduction of decoupled BLD, which is significantly more efficient than coupled BLD, makes it feasible to evaluate a much larger set of potential blocks within a single optimization run. This efficiency enables exploration of novel operations as alternative Puzzle blocks, such as variable window attention mechanisms [9], state-space models [18, 17], or other architectural innovations. The framework could also be extended to optimize models for specific capabilities, such as Chain-of-Thought reasoning or multimodal tasks, including vision-language models [38, 62] and retrieval-augmented generation [33]. [38] and retrieval-augmented generation [33].